OpenStudy (anonymous):

solve each system analytically. if the solution is dependent write solution set in terms of the variable z. x+y+z=2 2x+y-z=5 x-y+z=-2 I keep getting lost

OpenStudy (anonymous):

Have you learned how to solve systems of equations using matrices?

OpenStudy (anonymous):

I do not believe so. what is a matrice?

OpenStudy (anonymous):

A matrix is a way of representing a bunch of numbers. You don't need to worry about it if you haven't heard of it

OpenStudy (anonymous):

I would use the elimination method for systems of equations to start

OpenStudy (anonymous):

x=1 z=-1 y=2

OpenStudy (anonymous):

@JULIANCHICO it is against the rules to provide the answer straight out. From the code of conduct: "Don't just provide the answer to a problem when someone else is in the middle of helping! But if you want to help, by all means, join in!"

OpenStudy (anonymous):

Because the first equation and last equation both have positive x and positive z, I would combine those by multiplying one by negative one first: equation 1, multiplied by negative one: -x-y-z=-2 equation 3: x-y+z=-2 You can then combine them: (-x-y-z=-2) + (x-y+z=-2) ------------- -2y=-4

OpenStudy (anonymous):

That can get you strated @angelabink

OpenStudy (anonymous):

my bad @cholo71796

OpenStudy (anonymous):

ok that is a good start I will try from there

OpenStudy (anonymous):

ty